# Unraveling Open Quantum Dynamics with Time-Dependent Variational Monte Carlo

### v1

@@ Abstract
We introduce a method to simulate open quantum many-body dynamics by combining time-dependent variational Monte Carlo (tVMC) with quantum trajectory techniques. Our approach unravels the Lindblad master equation into an ensemble of stochastic Schrödinger equations for a variational ansatz, avoiding the exponential cost of density matrix evolution. The method is compatible with generic ansatzë, including expressive neural-network wavefunctions. We derive the nonlinear stochastic equations of motion for the variational parameters and employ suitable Stratonovich numerical solvers. To validate our approach, we simulate quenches in the locally dissipative long-range Ising model in a transverse field, accurately capturing non-equilibrium magnetization and spin squeezing dynamics relevant to trapped-ion and Rydberg atom experiments. The framework is computationally efficient, scalable on high-performance computing platforms, and can be readily integrated into existing tVMC implementations. This work enables large-scale simulations of complex, dissipative quantum systems in higher dimensions, with broad implications for quantum technology and fundamental science.
@@
Understanding and controlling the dynamics of open quantum many-body systems  [1, 2] is central to the development of quantum technologies. The roles of dissipation and decoherence have been extensively investigated across a range of areas, including quantum computation [3], quantum metrology [4, 5], quantum communication [6] and quantum thermodynamics [7]. In fact, the interplay of environment-induced noise with many-body coherence and correlations in driven, out-of-equilibrium regimes offers a fertile ground for investigating fundamental aspects of quantum physics, with direct implications for the study of thermalization, quantum coherence, entanglement, dissipative phase transitions and information flow in noisy environments [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. These investigations also pave the way for technological applications, particularly in the emerging domain of critical quantum metrology [19, 20, 21, 22, 23, 24, 25, 26]. A wide range of platforms—such as trapped ions, Rydberg atoms, ultracold gases, and superconducting circuits—has made the controlled realization of open quantum dynamics experimentally accessible, prompting the development of powerful theoretical tools.
@@
Due to the high dimensionality of the density matrix for quantum many-body systems, various approaches and numerical techniques have been devised to simulate Markovian master equations [27], which are particularly efficient in the semiclassical regime [28, 29] or employ variational techniques [30], and neural-network representations of the density matrix [31, 32, 33], trained via stochastic minimization of the Lindbladian residual. Matrix product state techniques are also an intense field of study and are accurate in the absence of entanglement volume law [34, 35, 36]. Complementarily, open libraries are available for exact benchmarks with relatively small systems [37, 38, 39, 40].
@@
In this work, we introduce a scalable and flexible variational scheme for simulating open quantum many-body dynamics described by a Markovian master equation in the Lindblad form, by combining the time-dependent variational Monte Carlo (tVMC) framework [41, 42, 43, 44, 45] with the quantum trajectory formalism [46, 47]. This hybrid approach reformulates the Lindblad evolution as an ensemble of stochastic pure-state trajectories, each evolved with a variationally optimal dynamics. It allows us to retain the expressive power and scalability of tVMC while incorporating dissipation nonperturbatively and without uncontrolled approximations. We apply this method to the long-range Ising model in a transverse field under local dissipation [48, 49, 50], a paradigmatic setting for investigating entanglement generation. We analyze the emergence of metrologically useful spin squeezing [51, 52, 53, 54, 4, 55] in regimes where strong correlations, long-range interactions [56], and noise coexist, highlighting the promise of variational trajectory-based approaches for simulating complex open-system dynamics beyond the reach of existing techniques.
@@
If the initial density matrix is given by $\hat{\rho}(0)=\sum_{i}p_{i}|\psi_{0,i}\rangle\langle\psi_{0,i}|$, where $p_{i}$ are classical probabilities and $|\psi_{0,i}\rangle$ are normalized pure states, then each quantum trajectory is initialized in the state $|\psi_{0,i}\rangle$ with probability $p_{i}$. The subsequent evolution of each trajectory follows the SSE (2). After generating $N_{\text{T}}$ such trajectories, denoted by $\{|\psi_{j}\rangle\}_{j=1,\dots,N_{T}}$, the expectation value of the observable $\hat{A}$ is obtained by ensemble averaging,
$\langle\!\langle\hat{A}\rangle\!\rangle=\frac{1}{N_{T}}\sum_{j}\langle\psi_{j}|\hat{A}|\psi_{j}\rangle\,\text{.}$ (3)
This trajectory-based calculation has a computational complexity of $O(N_{T}d)$. For example, choosing $N_{T}\approx 1000$ (which typically reduces the statistical error to a few percent) yields a complexity significantly lower than $O(d^{2})$, even for systems of moderate size. Furthermore, since each trajectory is independent, the calculation is trivially parallelizable, further reducing the overall computation time.
Despite these advantages, the Hilbert space dimension $d$ still grows exponentially with the system size. To address this difficulty, we approximate the quantum state of each trajectory using a variational ansatz of the form $e^{\phi}|\psi_{\bm{\theta}}\rangle$, where $\phi\in\mathbb{C}$ is a complex parameter controlling the global phase and norm, and $\bm{\theta}=(\theta_{1},\dots,\theta_{M})\in\mathbb{C}^{M}$ is a set of $M$ complex variational parameters. Given a configuration $x$ of the system, the corresponding variational wavefunction is expressed as $e^{\phi}\psi_{\bm{\theta}}(x)=e^{\phi}\langle x|\psi_{\bm{\theta}}\rangle$. This approach reduces the effective dimensionality of the state space from $d$ to $M+1$, lowering the computational complexity to $O(N_{T}M)$. The central task then becomes determining how the variational parameters evolve in time along each quantum trajectory.
@@
for $k=1,\dots,M$, where, by introducing the local operators $\mathcal{O}_{k}(x)=\frac{\partial}{\partial\theta_{k}}\ln\langle x|\psi_{\bm{\theta}}\rangle$, the local effective energy $\mathcal{E}_{\text{eff}}(x)=\langle x|\hat{H}^{\text{Strat}}_{\text{eff}}|\psi_{\bm{\theta}}\rangle/\langle x|\psi_{\bm{\theta}}\rangle$, and the $\mathcal{B}^{n}(x)=\langle x|\hat{L}_{n}|\psi_{\bm{\theta}}\rangle/\langle x|\psi_{\bm{\theta}}\rangle$, the quantum geometric tensor $S$ and the quantum force $F$ and noise $N^{n}$ vectors are defined as follows: $S_{k,k^{\prime}}=\langle\mathcal{O}_{k}^{*}\mathcal{O}_{k^{\prime}}\rangle-\langle\mathcal{O}_{k}^{*}\rangle\langle\mathcal{O}_{k^{\prime}}\rangle$, $F_{k}=\langle\mathcal{O}_{k}^{*}\mathcal{E}_{\text{eff}}\rangle-\langle\mathcal{O}_{k}^{*}\rangle\langle\mathcal{E}_{\text{eff}}\rangle$, and $N_{k}^{n}=\langle\mathcal{O}_{k}^{*}\mathcal{B}^{n}\rangle-\langle\mathcal{O}_{k}^{*}\rangle\langle\mathcal{B}^{n}\rangle$. With a common abuse of notation, here we indicate the Monte Carlo average of a local variable $\mathcal{A}$ with $\langle\mathcal{A}\rangle$.
@@
that depends on the solutions $d\theta_{k}$ of Eq. (8). Although this equation can be integrated to track the evolution of $\phi$, it is not necessary to do so in practice, since $\phi$ encodes only gauge degrees of freedom—the global phase and norm—which do not affect the physical information encoded in the state vector. Nevertheless, it plays an essential auxiliary role in the derivation of the variational equations of motion.
The system of equations (8) governs the evolution of the variational parameters $\bm{\theta}$ and lies at the core of our method, which we refer to as the open time-dependent variational Monte Carlo (otVMC). An otVMC simulation proceeds as follows. As is commonplace in the quantum diffusion approach, we assume a pure initial state. Within our variational representation, it is identified by a parameter vector $\bm{\theta}_{0}$, such that the initial density matrix is $\hat{\rho}(0)\propto|\psi_{\bm{\theta}_{0}}\rangle\langle\psi_{\bm{\theta}_{0}}|$. We then generate a set of $N_{T}$ quantum trajectories, each initialized from the state $|\psi_{\bm{\theta}_{0}}\rangle$.
@@
Figure 1: Evolution of magnetization components (panel a) and spin squeezing parameter (panel b) for $N=10$, $\kappa=0.5J$ and $h=0$. Comparison of different ansatzë (solid lines) to the exact result of App. D (dashed line). In panel a of all figures, errorbars are smaller than symbol sizes.
@@
We investigate the accuracy and computational cost of various variational ansatzë: a very expressive neural network quantum state of restricted Boltzmann machine (RBM) type [43]; and a less expensive long-range Jastrow (LJ) wavefunction with single and two-body correlations with different levels of non-uniformity. The RBM state is defined by:
@@
In Fig. 1, we show the time evolution of the components of magnetization (panel a) and the spin-squeezing parameter (panel b) for $N=10$ spins, $\kappa=0.5J$ and no transverse field. We benchmark the otVMC results with different ansatzë to the exact result from Ref. [63] (see App. D). All results are obtained by averaging over $N_{\text{T}}=400$ trajectories. We notice that all employed wavefunctions yield accurate results for short times, in particular for the magnetization along $z$. This is probably a merit of the inhomogeneous bias parameters. The RBM and LJNH wavefunctions are highly consistent with each other and are very accurate even for the spin-squeezing parameter around its minimum. The computationally cheaper LJPHd ansatz performs very well for magnetization and well for the challenging $\xi^{2}$ around its minimum, provided that $d>0$. In particular, while LJH does not sufficiently accommodate inhomogeneity, we argue that LJPH1 strikes a good compromise between scalability and accuracy.
@@
Figure 3: Study of magnetization and spin squeezing for $N=100$ and $\kappa/J=0.27$ with varying transverse field $h=0,-1.1,0.5$. We employ the LJPH1 ansatz (solid lines) and report the exact result of App. D for $h=0$ (dashed lines).
@@
Conclusions — We have shown an efficient and scalable time-dependent variational Monte Carlo method to simulate the Lindblad dynamics of many-body open quantum systems. Our approach leverages the unraveling of the master equation into stochastic Schrödinger trajectories within a compact yet expressive variational manifold, exemplified here by different ansatzë. We have focused on a one-dimensional long-range model which is highly non-trivial and relevant to experiments, considering the case without a transverse magnetic field, for benchmarking purposes, and the transverse field case, to demonstrate the flexibility of the method. Remarkably, our approach can scale to hundreds of particles, owing to its natural parallelizability, and is well suited to integration into existing codebases. Our method can be applied to more complex models in higher dimensions, where tensor network approaches are not guaranteed to be accurate in general. Other relevant research directions are the extension to dissipative systems in the continuum, like ultracold gases and liquids, and to quantum optics systems such as cavity-coupled atoms. Suitable sampling will also give access to mixed initial states, such as thermal ones [70][11]. Our solution to the dynamics of nonlinear trajectories will also provide a route to model dissipative systems under continuous measurement [71, 67].
Data to reproduce the figures in this paper are available online (Ref. [72]).
@@ Acknowledgements.
We thank T. Comparin for discussing with us the idea that fostered this work. We acknowledge stimulating discussions with T. Roscilde and G. Carleo. We acknowledge the early contributions of L. D’Onofrio and S. Lonardoni, developed in the framework of their theses. We acknowledge funding from Italian Ministry of Research and Next Generation EU via the PRIN 2022 Project CONTRABASS (Contract No. 2022KB2JJM). We acknowledge the CINECA awards IsCb5-PSIOQUAS and IsCc7-SDOQS under the ISCRA initiative, for the availability of high-performance computing resources and support. We acknowledge computational resources provided by INDACO Platform, which is a project of high performance computing at Università degli Studi di Milano. Christian Apostoli and Jacopo D’Alberto contributed equally to this work.
@@ Appendix D Exact results for the 1D LITF model
For the LITF without transverse field, we employ the analytical solution from [63] and take advantage of the translational invariance of the coupling, defining $J_{d}=(J/\mathcal{K}(\alpha))|d|^{-\alpha}$ for $d<N/2$ and $J_{d}=J_{N-d}$ for $d\geq N/2$. We consider a quench starting from the coherent spin state along $+x$, set $\Gamma=\kappa/2$, and define the following auxiliary quantities:

### v2

@@ Abstract
We introduce a method to simulate open quantum many-body dynamics by combining time-dependent variational Monte Carlo (tVMC) with quantum trajectory techniques. Our approach unravels the Lindblad master equation into an ensemble of stochastic Schrödinger equations for a variational ansatz, avoiding the exponential cost of density matrix evolution. The method is compatible with generic ansätze, including expressive neural-network wavefunctions. We derive the nonlinear stochastic equations of motion for the variational parameters and employ suitable Stratonovich numerical solvers. To validate our approach, we simulate quenches in the locally dissipative long-range Ising model in a transverse field, accurately capturing non-equilibrium magnetization and spin squeezing dynamics relevant to trapped-ion and Rydberg atom experiments. The framework is computationally efficient, scalable on high-performance computing platforms, and can be readily integrated into existing tVMC implementations. This work enables large-scale simulations of complex, dissipative quantum systems in higher dimensions, with broad implications for quantum technology and fundamental science.
@@
Understanding and controlling the dynamics of open quantum many-body systems  [1, 2] is central to the development of quantum technologies. The roles of dissipation and decoherence have been extensively investigated across a range of areas, including quantum computation [3], quantum metrology [4, 5], quantum communication [6] and quantum thermodynamics [7]. In fact, the interplay of environment-induced noise with many-body coherence and correlations in driven, out-of-equilibrium regimes offers a fertile ground for investigating fundamental aspects of quantum physics, with direct implications for the study of thermalization, quantum coherence, entanglement, dissipative phase transitions and information flow in noisy environments [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. These investigations also pave the way for technological applications, particularly in the emerging domain of critical quantum metrology [19, 20, 21, 22, 23, 24, 25, 26]. A wide range of platforms, such as trapped ions, Rydberg atoms, ultracold gases, and superconducting circuits, has made the controlled realization of open quantum dynamics experimentally accessible, prompting the development of powerful theoretical tools.
@@
Due to the high dimensionality of the density matrix for quantum many-body systems, various approaches and numerical techniques have been devised to simulate Markovian master equations [27], which are particularly efficient in the semiclassical regime [28, 29] or employ variational techniques [30], and neural-network representations of the density matrix [31, 32, 33], trained via stochastic minimization of the Lindbladian residual. Matrix product state techniques are also an intense field of study and are accurate in the absence of entanglement volume law [34, 35, 36, 37]. Complementarily, open libraries are available for exact benchmarks with relatively small systems [38, 39, 40, 41].
@@
In this work, we introduce a scalable and flexible variational scheme for simulating open quantum many-body dynamics described by a Markovian master equation in the Lindblad form, by combining the time-dependent variational Monte Carlo (tVMC) framework [42, 43, 44, 45, 46] with the quantum trajectory formalism [47, 48]. This hybrid approach reformulates the Lindblad evolution as an ensemble of stochastic pure-state trajectories, each evolved with a variationally optimal dynamics. It allows us to retain the expressive power and scalability of tVMC while incorporating dissipation nonperturbatively and without uncontrolled approximations. Access to the full time evolution and not only to the stationary state is crucial for quantum technologies, in particular for quantum state engineering and quantum metrology protocols. In this respect, we apply our method to the long-range Ising model in a transverse field under local dissipation [49, 50, 51], a paradigmatic setting for investigating entanglement generation. We analyze the emergence of metrologically useful spin squeezing during quench dynamics [52, 53, 54, 55, 4, 56] in regimes where strong correlations, long-range interactions [57], and noise coexist, highlighting the promise of variational trajectory-based approaches for simulating complex open-system dynamics beyond the reach of existing techniques.
@@
If the initial density matrix is given by $\hat{\rho}(0)=\sum_{i}p_{i}|\psi_{0,i}\rangle\langle\psi_{0,i}|$, where $p_{i}$ are classical probabilities and $|\psi_{0,i}\rangle$ are normalized pure states, then each quantum trajectory is initialized in the state $|\psi_{0,i}\rangle$ with probability $p_{i}$. The subsequent evolution of each trajectory follows the SSE (2). After generating $N_{\text{T}}$ such trajectories, denoted by $\{|\psi_{j}\rangle\}_{j=1,\dots,N_{\text{T}}}$, the expectation value of the observable $\hat{A}$ is obtained by ensemble averaging,
$\langle\!\langle\hat{A}\rangle\!\rangle=\frac{1}{N_{\text{T}}}\sum_{j}\langle\psi_{j}|\hat{A}|\psi_{j}\rangle\,\text{.}$ (3)
This trajectory-based calculation has a computational complexity of $O(N_{\text{T}}d)$. For example, choosing $N_{\text{T}}\approx 1000$ (which typically reduces the statistical error to a few percent) yields a complexity significantly lower than $O(d^{2})$, even for systems of moderate size. Furthermore, since each trajectory is independent, the calculation is trivially parallelizable, further reducing the overall computation time.
Despite these advantages, the Hilbert space dimension $d$ still grows exponentially with the system size. To address this difficulty, we approximate the quantum state of each trajectory using a variational ansatz of the form $e^{\phi}|\psi_{\bm{\theta}}\rangle$, where $\phi\in\mathbb{C}$ is a complex parameter controlling the global phase and norm, and $\bm{\theta}=(\theta_{1},\dots,\theta_{M})\in\mathbb{C}^{M}$ is a set of $M$ complex variational parameters. Given a configuration $x$ of the system, the corresponding variational wavefunction is expressed as $e^{\phi}\psi_{\bm{\theta}}(x)=e^{\phi}\langle x|\psi_{\bm{\theta}}\rangle$. This approach reduces the effective dimensionality of the state space from $d$ to $M+1$, lowering the computational complexity to $O(N_{\text{T}}M)$. The central task then becomes determining how the variational parameters evolve in time along each quantum trajectory.
@@
for $k=1,\dots,M$, where, by introducing the local operators $\mathcal{O}_{k}(x)=\frac{\partial}{\partial\theta_{k}}\ln\langle x|\psi_{\bm{\theta}}\rangle$, the local effective energy $\mathcal{E}_{\text{eff}}(x)=\langle x|\hat{H}^{\text{Strat}}_{\text{eff}}|\psi_{\bm{\theta}}\rangle/\langle x|\psi_{\bm{\theta}}\rangle$, and the local noises $\mathcal{B}^{n}(x)=\langle x|\hat{L}_{n}|\psi_{\bm{\theta}}\rangle/\langle x|\psi_{\bm{\theta}}\rangle$, the quantum geometric tensor $S$ and the quantum force $F$ and noise $N^{n}$ vectors are defined as follows: $S_{k,k^{\prime}}=\langle\mathcal{O}_{k}^{*}\mathcal{O}_{k^{\prime}}\rangle-\langle\mathcal{O}_{k}^{*}\rangle\langle\mathcal{O}_{k^{\prime}}\rangle$, $F_{k}=\langle\mathcal{O}_{k}^{*}\mathcal{E}_{\text{eff}}\rangle-\langle\mathcal{O}_{k}^{*}\rangle\langle\mathcal{E}_{\text{eff}}\rangle$, and $N_{k}^{n}=\langle\mathcal{O}_{k}^{*}\mathcal{B}^{n}\rangle-\langle\mathcal{O}_{k}^{*}\rangle\langle\mathcal{B}^{n}\rangle$. With a common abuse of notation, here we indicate the Monte Carlo average of a local variable $\mathcal{A}$ with $\langle\mathcal{A}\rangle$.
@@
that depends on the solutions $d\theta_{k}$ of Eq. (8). Although this equation can be integrated to track the evolution of $\phi$, it is not necessary to do so in practice, since $\phi$ encodes only gauge degrees of freedom, namely, the global phase and norm, which do not affect the physical information encoded in the state vector. Nevertheless, it plays an essential auxiliary role in the derivation of the variational equations of motion.
The system of equations (8) governs the evolution of the variational parameters $\bm{\theta}$ and lies at the core of our method, which we refer to as the open time-dependent variational Monte Carlo (otVMC). An otVMC simulation proceeds as follows. As is commonplace in the quantum diffusion approach, we assume a pure initial state. Within our variational representation, it is identified by a parameter vector $\bm{\theta}_{0}$, such that the initial density matrix is $\hat{\rho}(0)\propto|\psi_{\bm{\theta}_{0}}\rangle\langle\psi_{\bm{\theta}_{0}}|$. We then generate a set of $N_{\text{T}}$ quantum trajectories, each initialized from the state $|\psi_{\bm{\theta}_{0}}\rangle$.
@@
Figure 1: Evolution of magnetization components (panel a) and spin squeezing parameter (panel b) for $N=10$, $\kappa=0.5J$ and $h=0$. Comparison of different ansätze (solid lines) to the exact result of App. D (dashed line). In panel a of all figures, uncertainty bars are smaller than symbol sizes.
We investigate the accuracy and computational cost of various variational ansätze: a very expressive neural network quantum state of restricted Boltzmann machine (RBM) type [44]; and a less expensive long-range Jastrow (LJ) wavefunction with single and two-body correlations with different levels of non-uniformity. The RBM state is defined by:
@@
In Fig. 1, we show the time evolution of the components of magnetization (panel a) and the spin-squeezing parameter (panel b) for $N=10$ spins, $\kappa=0.5J$ and no transverse field. We benchmark the otVMC results with different ansätze to the exact result from Ref. [64] (see App. D). All results are obtained by averaging over $N_{\text{T}}=400$ trajectories. We notice that all employed wavefunctions yield accurate results for short times, in particular for the magnetization along $z$. This is probably a merit of the inhomogeneous bias parameters. The RBM and LJNH wavefunctions are highly consistent with each other and are very accurate even for the spin-squeezing parameter around its minimum. The computationally cheaper LJPHd ansatz performs very well for magnetization and well for the challenging $\xi^{2}$ around its minimum, provided that $d>0$. In particular, while LJH does not sufficiently accommodate inhomogeneity, we argue that LJPH1 strikes a good compromise between scalability and accuracy. See App. E for considerations on the computational cost of our simulations.
@@
Conclusions — We have shown an efficient and scalable time-dependent variational Monte Carlo method to simulate the Lindblad dynamics of many-body open quantum systems. Our approach leverages the unraveling of the master equation into stochastic Schrödinger trajectories within a compact yet expressive variational manifold, exemplified here by different ansätze.
Figure 3: Study of magnetization and spin squeezing for $N=100$ and $\kappa/J=0.27$ with varying transverse field $h=0,-0.53,1.07$. We employ the LJPH1 and LJH ansätze (solid lines) and report the exact result of App. D for $h=0$ (dashed lines).
We have focused on a one-dimensional long-range model which is highly non-trivial and relevant to experiments, considering the case without a transverse magnetic field, for benchmarking purposes, and the transverse field case, to demonstrate the flexibility of the method. Remarkably, our approach can scale to hundreds of particles, owing to its natural parallelizability, and is well suited to integration into existing codebases. It can access not only the stationary state, but the full open quantum dynamics, which is important for modeling quantum technologies. Our method can be applied to more complex models in higher dimensions, where tensor network approaches are not guaranteed to be accurate in general. Other relevant research directions are the extension to dissipative systems in the continuum, like ultracold gases and liquids, to quantum optics systems such as cavity-coupled atoms, and to the study of defect formation in quenches [62]. Suitable sampling will also give access to mixed initial states, such as thermal ones [71, 11]. Our solution to the dynamics of nonlinear trajectories will also provide a route to model dissipative systems under continuous measurement [72, 68].
The data that support the findings of this article are openly available [73].
@@ Acknowledgements.
We thank T. Comparin for discussing with us the idea that fostered this work. We acknowledge stimulating discussions with T. Roscilde and G. Carleo. We acknowledge the early contributions of L. D’Onofrio and S. Lonardoni, developed in the framework of their theses. We acknowledge funding from Italian Ministry of Research and Next Generation EU via the PRIN 2022 Project CONTRABASS (Contract No. 2022KB2JJM). We acknowledge the CINECA awards IsCb5-PSIOQUAS and IsCc7-SDOQS under the ISCRA initiative, for the availability of high-performance computing resources and support. We acknowledge computational resources provided by INDACO Platform, which is a project of high performance computing at Università degli Studi di Milano.
@@ Appendix D Exact results for the 1D LITF model
For the LITF without transverse field, we employ the analytical solution from [64] and take advantage of the translational invariance of the coupling, defining $J_{d}=(J/\mathcal{K}(\alpha))|d|^{-\alpha}$ for $20<d<N/$ 2, $J_{d}=J_{N-d}$ for $d\geq N/2$ and $J_{0}=0$. We consider a quench starting from the coherent spin state along $+x$, set $\Gamma=\kappa/2$, and define the following auxiliary quantities:
@@
Appendix E Computational cost
N LJH LJPH1 LJPH2 LJNH RBM
10 7s 9s 13s 25s 2min
100 15min
160 30min
Table 1: Wall time duration of typical trajectories for total time $T_{\text{max}}\approx 0.8/J$. See text.
In our method, both tVMC sampling and the quantum trajectories can be parallelized, and our code currently employs MPI on multiple CPUs. We collect in Table 1 the typical wall times for single quantum trajectories simulated in this work, with tVMC sampling distributed on 72 cores @2.5GHz. The LJPH1 ansatz simulations scale approximately as $N^{2}$.